OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{\normalsize \text{A can run around a circular track in 4 min.} \hspace{.33em}\\~\\ \normalsize \text{ B can do the same thing in 7 min. } \hspace{.33em}\\~\\ \normalsize \text{ If they start running together from same point , clock wise } \hspace{.33em}\\~\\ \normalsize \text{when will they meet for the first time } \hspace{.33em}\\~\\ \normalsize \text{ at the point diametrically opposite to} \hspace{.33em}\\~\\ \normalsize \text{ the starting point. ?}\hspace{.33em}\\~\\}\end{align}\)

OpenStudy (mathmath333):



OpenStudy (zarkon):

if their speeds are constant then, unless i'm missing something, they will never meet at the same time

OpenStudy (zarkon):

A will be at the opposite side at 2,6,10,... B will be at the opposite side at 3.5,11.5,17.5,... (always on a half min)

OpenStudy (zarkon):

3.5,10.5,17.5...

OpenStudy (mathmath333):

yes zarkon is right!

OpenStudy (mathmath333):

since the distance is not given we can also assume it

OpenStudy (mathmath333):

 \(\large \begin{align} \color{black}{t_1=\dfrac{d}{s_1} \hspace{.33em}\\~\\ =\dfrac{14}{7} \hspace{.33em}\\~\\ =2,6,10,14..... \hspace{.33em}\\~\\ t_2=\dfrac{14}{s_2} \hspace{.33em}\\~\\ =\dfrac{14}{2} \hspace{.33em}\\~\\ =7,21,35,49..... \hspace{.33em}\\~\\ }\end{align}\)